The position of particle is given as =3-4t+3t2m,what is the instantaneous acceleration at t=2s


1
Expert's answer
2021-02-22T11:07:03-0500

Position of particle as function of time is given as

"s(t)=3-4t+3t^2" m

Differentiating above equation with respect to time,

"\dfrac{ds}{dt}=-4+6t"

Again differentiating the above equation with respect to time,

"\dfrac{d^2s}{dt^2}=6"

"a=6ms^{-2}"

Since acceleration of the particle is independent of time, therefore

instantaneous acceleration at time t=2s will be

"a=6ms^{-2}"